OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? -3x+4y=-20 x+4y=-4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (caozeyuan):

if x=a, y=b, then the ordered pair(a,b) is the solution of the system if the LHS equals the RHS

OpenStudy (anonymous):

im really bad at math

OpenStudy (caozeyuan):

less technical saying is that take you first number of your ordered pair, substitue all x with this number, do the smae thing on y with the second number, now compute your left side of equations, if it equals the right side, you found the solution

OpenStudy (anonymous):

huh?

OpenStudy (caozeyuan):

so your first choice is (8,1) which means you should go change every x to 8 and every y to 1

OpenStudy (anonymous):

mk

OpenStudy (caozeyuan):

now what's your equation look like?

OpenStudy (caozeyuan):

so the LHS of yor first equation used to be -3x+4, now its -3*8+4 which is -20

OpenStudy (anonymous):

-8x+1y=-20 8x+1y=-4

OpenStudy (caozeyuan):

-3x+4y, sorry

OpenStudy (caozeyuan):

nonono

OpenStudy (caozeyuan):

change your x's and y's not your constants

OpenStudy (anonymous):

-8x+1y

OpenStudy (caozeyuan):

why do you do that?!

OpenStudy (caozeyuan):

x=8 and y=1, so if you have x, change it to 8, you have y change to 1

OpenStudy (anonymous):

sry -8+1=

OpenStudy (caozeyuan):

did you figure out how to do it ?

OpenStudy (anonymous):

no i dont understand anything to do with math besides adding subtracting and multiplication and division everything else i suck at.

OpenStudy (caozeyuan):

ok, start from basics, what does (8,1) stand for?

OpenStudy (anonymous):

idk

OpenStudy (caozeyuan):

it means x=8 and y=1

OpenStudy (anonymous):

the x and y axis?

OpenStudy (caozeyuan):

(a,b) means x=a and y=b

OpenStudy (anonymous):

ok

OpenStudy (caozeyuan):

geometircally yes, but you can just say algebraically, without talking about the coordinate plane which makes things easier cuz you are not doing geometry now

OpenStudy (anonymous):

i kno im doing pre-algibra

OpenStudy (caozeyuan):

next, what it the solution to a equation or a system of equations

OpenStudy (caozeyuan):

what?

OpenStudy (caozeyuan):

they teach algebra in pre-algebra, this doesnt een make sense

OpenStudy (anonymous):

i have a very slow brain

OpenStudy (anonymous):

so sometimes i dont make sense

OpenStudy (caozeyuan):

for an ordered pair (a,b) to be a system of equations, x=a and y=b MUST hold true for both equations

OpenStudy (caozeyuan):

in your example -3x+4y=-20, x+4y=-4

OpenStudy (caozeyuan):

say order pair (a,b) is the solution of equation, -3a+4b=20 AND a+4bb=-4 MUST BOTH be true

OpenStudy (caozeyuan):

therefore (1,1) is NOT a solution since -3*1+4 is not -20

OpenStudy (caozeyuan):

understand better?

OpenStudy (anonymous):

im sorry no

OpenStudy (caozeyuan):

what are you failing at? now I tried my best, if you still dont understand, you need to tell me where EXACTLY you are not understanding

OpenStudy (anonymous):

ok so im not understanding any of it i kno you try you best you might as well just quit because i cant understand math

OpenStudy (caozeyuan):

one last round, ok

OpenStudy (anonymous):

go for it

OpenStudy (caozeyuan):



OpenStudy (caozeyuan):

so far so goo?

OpenStudy (anonymous):

okay

OpenStudy (caozeyuan):